# Centered Moments of Weighted One-Level Densities of $GL(2)$ $L$-Functions

### v3

@@ 2. Notation and Preliminaries
We follow the definitions and notations in [IK04] throughout. We define the $N$th congruence subgroup of $SL_{2}(\mathbb{Z})$ by
$\Gamma_{0}(N)\ :=\ \left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL_{2}(\mathbb{Z})\ \Bigg|\ c\equiv 0\pmod{N}\right\}$. (2.1)
@@
$f\left(\frac{az+b}{cz+d}\right)=(cz+d)^{k}f(z).$ (2.2)
@@
$\langle f,g\rangle\ :=\ \frac{1}{\nu(N)}\int_{\Gamma_{0}(N)\backslash\mathbb{H}}f(z)\overline{g(z)}\,y^{k}\,\frac{dx\,dy}{y^{2}},$ (2.3)
@@
$T_{n}f(z)\ :=\ n^{k-1}\sum_{\begin{subarray}{c}ad=n\\ (a,N)=1\end{subarray}}\sum_{b=0}^{d-1}d^{-k}f\left(\frac{az+b}{d}\right).$ (2.4)
The theory of functional analysis provides that the Hecke operators are simultaneously diagonalizable, allowing us to speak of an orthogonal basis $\mathcal{F}_{k}(N)$ of forms which are simultaneous eigenforms of all these Hecke operators (Hecke eigenforms for short). The forms $f\in\mathcal{F}_{k}(N)$ admit a Fourier expansion
$f(z)=\sum_{n=1}^{\infty}a_{f}(n)e^{2\pi inz},$ (2.5)
@@
$\left|\mathcal{F}_{k}(N)\right|\sim\frac{k-1}{12}\phi(N)+O\left((kN)^{5/6}\right),$ (2.6)
where $\phi$ is Euler’s totient function. For any Hecke eigenform $f\in\mathcal{F}_{k}(N)$, we refer to the eigenvalue of $f$ under $T_{n}$ as the $n$th Hecke eigenvalue of $f$. It is straightforward from Definitions 2.5 and 2.4 that the Hecke eigenvalues of $f$ are closely related to the Fourier coefficients of $f$:
$a_{f}(n)\ =\ \lambda_{f}(n)n^{(k-1)/2}.$ (2.7)
Furthermore, the Hecke eigenvalues of $f$ possess useful multiplicative properties:
$\lambda_{f}(m)\lambda_{f}(n)\ =\ \sum_{\begin{subarray}{c}d\mid(m,n)\\ (d,N)=1\end{subarray}}\lambda_{f}\left(\frac{mn}{d^{2}}\right);$ (2.8)
@@
$\lambda_{f}(m)\lambda_{f}(n)\ =\ \lambda_{f}(mn),$ (2.9)
and if $p$ is a prime not dividing $N$, then
$\lambda_{f}(p)^{2k}\$$=\ \sum_{r=0}^{k}\left(\binom{2k}{k-r}-\binom{2k}{k-r-1}\right)\lambda_{f}(p^{2k}),$ (2.10)
$\lambda_{f}(p)^{2k+1}\$$=\ \sum_{r=0}^{k}\left(\binom{2k+1}{k-r}-\binom{2k+1}{k-r-1}\right)\lambda_{f}(p^{2k+1}).$ (2.11)
Proofs of (2.10) and (2.11) can be found in [Guy00]. These properties motivate the definition of the $L$-function associated to $f$:
@@
$L(s,f\times\chi)\ :=\ \sum_{n=1}^{\infty}\frac{\chi(n)\lambda_{f}(n)}{n^{s}},\quad\Re(s)>1.$ (2.13)
@@
$\Lambda(s,f\times\chi)\ :=\ (2\pi)^{-s-\frac{k-1}{2}}\Gamma\left(s+\frac{k-1}{2}\right)L(s,f\times\chi).$ (2.14)
@@
$\Lambda(s,f\times\chi)\ =\ \frac{i^{k}}{D^{2s-1}}\frac{\tau(\chi)^{2}}{D}\ \Lambda(1-s,f\times\overline{\chi}),$ (2.15)
@@
  • $\mathcal{F}_{2}=\bigcup_{k,N}\mathcal{F}_{k}(N)^{\prime}$ with $k>2$ ranging over even integers satisfying $\tau(\chi)^{2}=-i^{k}D$, equivalently $\chi(-1)=-i^{k}$, and $N$ ranging over primes not divisible by $rD$.
@@
$w_{\chi,r}(f)\ :=\ \frac{\Lambda\left(\frac{1}{2},f\times\chi\right)|a_{f}(r)|^{2}}{\lVert f\rVert^{2}},$ (1.12)
@@ 3. Explicit Formula
In this section, we restrict our attention to a single sub-family $\mathcal{F}_{k}(N)^{\prime}$ of either $\mathcal{F}_{1}$ or $\mathcal{F}_{2}$. Given any Hecke eigenform $f\in\mathcal{F}_{k}(N)^{\prime}$, an explicit formula for the one-level density of $L(s,f)$ is given in [ILS00, (4.18)]:
@@
$-2\sum_{p\nmid N}\lambda_{f}(p^{2})\widehat{\phi}\left(\frac{2\log p}{\log Q_{k,N}}\right)\frac{\log p}{p\log Q_{k,N}}.$ (3.1)
Using crude estimates for the higher symmetric power $L$-function $L(s,\text{sym}^{2}(f)\otimes\text{sym}^{2}(f))$, we can absorb the second moment of the Hecke eigenvalues (3.1) into the error term $O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$, as explained in [ILS00, Appendix B].
As stated before, Knightly and Reno studied the average weighted one-level density of $\mathcal{F}_{k}(N)^{\prime}$, proving that it depends on the triviality of the twisting character in the weight [KR19, (4.8)]:
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\ =\ \begin{cases}\widehat{\phi}(0)-\frac{1}{2}\phi(0)+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&\text{if }\chi\text{ trivial},\\ \widehat{\phi}(0)+\frac{1}{2}\phi(0)+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&\text{if }\chi\text{ non-trivial}.\end{cases}$ (3.2)
@@
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(D(\cdot,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\right)^{n}\right].$ (3.3)
Applying (3.1) and (3.2), we determine that (3.3) equals
$O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)+\begin{cases}&\sum_{t=0}^{n}\binom{n}{t}(-2)^{t}\ \mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum_{p\nmid N}\lambda_{\cdot}(p)\widehat{\phi}\left(\frac{\log p}{\log(k^{2}N)}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right]\\ &\text{if }\chi\text{ trivial},\\ &(-2)^{n}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum_{p\nmid N}\lambda_{\cdot}(p)\widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{n}\right]\\ &\text{if }\chi\text{ non-trivial}.\end{cases}$ (3.4)
@@
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum\nolimits_{p\nmid N}\lambda_{\cdot}(p)\ \widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right],$ (3.5)
@@
$\left.\times\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j},\\ m_{j}\equiv n_{j}(2)\end{subarray}}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\lambda_{\cdot}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right)\right],$ (3.6)
where each $c_{n_{j},m_{j}}$ is bounded by an absolute constant in $t$.
@@
$\sum_{\begin{subarray}{c}(p_{1},\dots,p_{t})\\ p_{i}\nmid N\end{subarray}}\left[\left(\prod_{i=1}^{t}\widehat{\phi}\left(\frac{\log p_{i}}{\log Q_{k,N}}\right)\frac{\log p_{i}}{\sqrt{p_{i}}\log Q_{k,N}}\right)\ \mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\prod\nolimits_{i=1}^{t}\lambda_{\cdot}(p_{i})\right)\right].$ (3.7)
@@
$\frac{1}{\ell!}\binom{t}{n_{1}}\binom{t-n_{1}}{n_{2}}\cdots\binom{n_{\ell}}{n_{\ell}}\ =\ \frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}.$ (3.8)
@@
$\left.\times\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\prod\nolimits_{j=1}^{\ell}\lambda_{\cdot}(q_{j})^{n_{j}}\right)\right].$ (3.9)
As in (2.10) and (2.11), we write $\lambda_{\cdot}(q_{j})^{n_{j}}$ as a linear combination of Hecke eigenvalues at the powers of $q_{j}$:
$\lambda_{\cdot}(q_{j})^{n_{j}}\ =\ \sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}(2)\end{subarray}}c_{n_{j},m_{j}}\ \lambda_{\cdot}(q_{j}^{m_{j}}),$ (3.10)
where each $c_{n_{j},m_{j}}$ is bounded by an absolute constant in $n_{j}\leq t$. Making this substitution in (3.9) gives (3.6). ∎
@@ 4. Weighted Trace Formula
In this section, we derive an asymptotic trace formula for the weighted average of the $m$th Hecke eigenvalue ($m$ arbitrary) over $\mathcal{F}_{k}(N)^{\prime}$:
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(\lambda_{\cdot}(m))\ :=\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$ (4.1)
Trace formulae are more readily attainable for $\mathcal{F}_{k}(N)$ than $\mathcal{F}_{k}(N)^{\prime}$. This is because the equidistribution laws that govern Hecke eigenvalues—and thereby influence trace formulae—are better understood for the space of all forms $S_{k}(N)$ rather than the space of newforms alone. Studying families of newforms usually entails extensive bookkeeping and adapting trace formulae from the ambient space to the space of newforms, e.g., [ILS00, Propositions 2.1-2.8]. Fortunately, we do not have to do any bookkeeping in our study, owing to the strategic conditioning on $k$ and $N$ in our definitions of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$.
@@
$\Lambda\left(\frac{1}{2},g_{N}\times\chi\right)\ =\ \frac{\chi(N)}{N^{k/2}}\Lambda\left(\frac{1}{2},g\times\chi\right)\ =\ 0\implies w_{\chi,r}(g_{N})=0.$ (4.2)
It follows that the weight vanishes for all forms in the span of $\{g,g_{N}\}$, $f$ in particular. This is to say that $w_{\chi,r}(f)$ vanishes for all oldforms $f\in\mathcal{F}_{k}(N)$; hence any $w_{\chi,r}$-weighed average over $\mathcal{F}_{k}(N)$ reduces to a $w_{\chi,r}$-weighted average over $\mathcal{F}_{k}(N)^{\prime}$, and we can refer to these interchangeably.
In the interest of finding a weighted trace formula for (4.1), we call attention to a related formula in [JK15, Theorem 1.1]. By applying a certain relative trace formula for $S_{k}(N)$, Jackson and Knightly derived a formula for the sum of Fourier coefficients over $\mathcal{F}_{k}(N)$ weighted by
$w_{\chi,r}^{\prime}(f)\ :=\ \frac{\Lambda\left(\frac{1}{2},f\times\chi\right)\overline{a_{f}(r)}}{\lVert f\rVert^{2}},\quad f\in\mathcal{F}_{k}(N).$ (4.3)
Knightly and Reno analyzed this to obtain an asymptotic formula [KR19, (3.2)]. To be explicit, if $k>2$ and $N=1$, then for all $m\geq 1$,
$\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)a_{f}(m)\$$=\ \frac{2^{k-1}(2\pi)^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)(rm)^{\frac{k}{2}-1}\chi(rm)\ \sigma_{1}((r,m))\$
$+\ O\left(\frac{(4\pi rm)^{k-1}D^{\frac{k}{2}-\frac{1}{2}}}{(k-2)!}\right);$ (4.4)
and if $k>2$ and $N>1$, then for all $m\geq 1$,
$\frac{1}{\nu(N)}\sum_{f\in\mathcal{F}_{k}(N)}w_{\chi,r}^{\prime}(f)a_{f}(m)\$$=\ \frac{2^{k-1}(2\pi)^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\ (rm)^{\frac{k}{2}-1}\chi(rm)\ \sigma_{1}((r,m))$
$+\ O\left(\frac{(4\pi rm)^{k-1}D^{\frac{k}{2}-\frac{1}{2}}}{N^{\frac{k}{2}}(k-2)!}\right),$ (4.5)
where the implied constants depend only on $r$ and $D$, and $\sigma_{1}$ is the divisor sum function. (For the interested reader, we have specialized [KR19, (3.2)] for trivial nebentypus, real $\chi$, and $s=1/2$, and absorbed constants in $r$ and $D$ into the implied constant.)
In the following three lemmas, we adapt these asymptotic trace formulae for the weighted average of our interest (4.1).
@@
Suppose $k>2$ and $N=1$. Then, for any positive integer $m\nmid N$, the sum of the $m$th Hecke eigenvalue over $\mathcal{F}_{k}(1)^{\prime}$ is equal to
$\sum_{f\in\mathcal{F}_{k}(1)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)\$$=\ \frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)\sigma_{1}(r)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)m^{-\frac{1}{2}}\chi(m)\ \sigma_{1}((r,m))$
$+\ O\left(\frac{m^{\frac{k-1}{2}}V^{k}}{(k-2)!}\right),$ (4.6)
where the implied constant and $V$ depend only on $r$ and $D$.

### v4

@@ 2. Notation and Preliminaries
We follow the definitions and notation in [IK04] throughout. We define the $N$th congruence subgroup of $SL_{2}(\mathbb{Z})$ by
(2.1) $\Gamma_{0}(N)\ :=\ \left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL_{2}(\mathbb{Z})\ \Bigg|\ c\equiv 0\pmod{N}\right\}$.
@@
(2.2) $f\left(\frac{az+b}{cz+d}\right)=(cz+d)^{k}f(z).$
@@
(2.3) $\langle f,g\rangle\ :=\ \frac{1}{\nu(N)}\int_{\Gamma_{0}(N)\backslash\mathbb{H}}f(z)\overline{g(z)}\,y^{k}\,\frac{dx\,dy}{y^{2}},$
@@
(2.4) $T_{n}f(z)\ :=\ n^{k-1}\sum_{\begin{subarray}{c}ad=n\\ (a,N)=1\end{subarray}}\sum_{b=0}^{d-1}d^{-k}f\left(\frac{az+b}{d}\right).$
For all $n$ satisfying $(n,N)=1$, the Hecke operators $T_{n}$ simultaneously diagonalize in $S_{k}(N)$, so we may speak of an orthogonal basis $\mathcal{F}_{k}(N)$ of forms which are simultaneous eigenforms of all these Hecke operators (Hecke eigenforms for short). The forms $f\in\mathcal{F}_{k}(N)$ admit a Fourier expansion
(2.5) $f(z)=\sum_{n=1}^{\infty}a_{f}(n)e^{2\pi inz},$
@@
(2.6) $\left|\mathcal{F}_{k}(N)\right|\sim\frac{k-1}{12}\phi(N)+O\left((kN)^{5/6}\right),$
where $\phi$ is Euler’s totient function. For any Hecke eigenform $f\in\mathcal{F}_{k}(N)$, we refer to the eigenvalue of $f$ under $T_{n}$ as the $n$th Hecke eigenvalue of $f$. It is straightforward from (2.4) and (2.5) that the Hecke eigenvalues of $f$ are closely related to the Fourier coefficients of $f$:
(2.7) $a_{f}(n)\ =\ \lambda_{f}(n)n^{(k-1)/2}.$
Furthermore, the Hecke eigenvalues of $f$ possess useful multiplicative properties: for all $m,n\in\mathbb{N}$,
(2.8) $\lambda_{f}(m)\lambda_{f}(n)\ =\ \sum_{\begin{subarray}{c}d\mid(m,n)\\ (d,N)=1\end{subarray}}\lambda_{f}\left(\frac{mn}{d^{2}}\right);$
@@
(2.9) $\lambda_{f}(m)\lambda_{f}(n)\ =\ \lambda_{f}(mn),$
and if $p$ is a prime not dividing $N$ and $\ell\geq 0$ is an integer, then
(2.10) $\lambda_{f}(p)^{\ell}\$$=\ \sum_{\begin{subarray}{c}0\leq r\leq\ell\\ r\equiv\ell(2)\end{subarray}}^{\ell}c_{\ell,r}\cdot\lambda_{f}(p^{r}),$
where $c_{\ell,r}$ is defined recursively for all $\ell^{\prime}\geq 1$ by
$c_{1,1}\$:= $:=\ 1,$
$c_{2\ell^{\prime},2r^{\prime}}\$$:=\ c_{2\ell^{\prime}-1,2r^{\prime}-1}+c_{2\ell^{\prime}-1,2r^{\prime}+1}$ $\text{ for all }0<r^{\prime}<\ell^{\prime},$
$c_{2\ell^{\prime},0}\$$:=\ c_{2\ell^{\prime}-1,1},$
$c_{2\ell^{\prime},2\ell^{\prime}}\$$:=\ c_{2\ell^{\prime}-1,2\ell^{\prime}-1},$
$c_{2\ell^{\prime}+1,2r^{\prime}+1}\$$:=c_{2\ell^{\prime},2r^{\prime}}+c_{2\ell^{\prime},2r^{\prime}+2}$ $\text{ for all }0\leq r^{\prime}<\ell^{\prime},$
(2.11) $c_{2\ell^{\prime}+1,2\ell^{\prime}+1}\$$:=c_{2\ell^{\prime},2\ell^{\prime}}$ $\text{ for all }0\leq r^{\prime}<\ell^{\prime}.$
Using Stirling’s formula, we have the following crude bound for $c_{\ell,r}$: for all $1\leq r\leq\ell$,
(2.12) $c_{\ell,r}\ \leq\ 2^{\ell}.$
The multiplicative properties of the Hecke eigenvalues of $f$ motivate the definition of the $L$-function associated to $f$:
@@
(2.14) $L(s,f\times\chi)\ :=\ \sum_{n=1}^{\infty}\frac{\chi(n)\lambda_{f}(n)}{n^{s}},\quad\Re(s)>1.$
@@
(2.15) $\Lambda(s,f\times\chi)\ :=\ (2\pi)^{-s-\frac{k-1}{2}}\Gamma\left(s+\frac{k-1}{2}\right)L(s,f\times\chi).$
@@
(2.16) $\Lambda(s,f\times\chi)\ =\ \frac{i^{k}}{D^{2s-1}}\frac{\tau(\chi)^{2}}{D}\ \Lambda(1-s,f\times\overline{\chi}),$
@@
  • $\mathcal{F}_{2}=\bigcup_{k,N}\mathcal{F}_{k}(N)^{\prime}$ with $k>2$ ranging over even integers satisfying $\tau(\chi)^{2}=-i^{k}D$, equivalently $\chi(-1)=-i^{k}$, and $N$ ranging over primes not divisible by $rD$. We order the sub-families $\mathcal{F}_{k}(N)$ so that $k+N$ is increasing.
@@
(2.19) $w_{\chi,r}(f)\ :=\ \frac{\Lambda\left(\frac{1}{2},f\times\chi\right)|a_{f}(r)|^{2}}{\lVert f\rVert^{2}},$
@@ 3. Explicit Formula
In this section, we restrict our attention to a single sub-family $\mathcal{F}_{k}(N)^{\prime}$ of either $\mathcal{F}_{1}$ or $\mathcal{F}_{2}$. Given any Hecke eigenform $f\in\mathcal{F}_{k}(N)^{\prime}$ and any test function $\phi$ whose Fourier transform $\widehat{\phi}$ is compactly supported, an explicit formula for the one-level density of $L(s,f)$ is given in [ILS00, (4.18)]:
@@
(3.1) $-2\sum_{p\nmid N}\lambda_{f}(p^{2})\widehat{\phi}\left(\frac{2\log p}{\log Q_{k,N}}\right)\frac{\log p}{p\log Q_{k,N}}.$
Assuming $\text{supp}(\widehat{\phi})\subset\left(-\frac{1}{2},\frac{1}{2}\right)$, we know from [ILS00, 88] that the second summation is
(3.2) $\sum_{p\nmid N}\lambda_{f}(p^{2})\widehat{\phi}\left(\frac{2\log p}{\log Q_{k,N}}\right)\frac{\log p}{p\log Q_{k,N}}\ \ll\ \frac{\log\log(kN)}{\log Q_{k,N}}.$
Furthermore, Knightly and Reno studied the average weighted one-level density of $\mathcal{F}_{k}(N)^{\prime}$ for $\phi$ restricted as above, proving that it depends on the triviality of the twisting character in the weight [KR19, 165]:
(3.3) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\ =\ \begin{cases}\widehat{\phi}(0)-\frac{1}{2}\phi(0)+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&\text{if }\chi\text{ trivial},\\ \widehat{\phi}(0)+\frac{1}{2}\phi(0)+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&\text{if }\chi\text{ non-trivial}.\end{cases}$
@@
(3.4) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(D(\cdot,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\right)^{n}\right].$
Subtracting (3.3) from (3.1), we determine that (3.4) equals
[…]
@@
(3.6) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum\nolimits_{p\nmid N}\lambda_{\cdot}(p)\ \widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right],$
@@
$\sum_{\begin{subarray}{c}1\leq\ell\leq t,\\ n_{1}+\dots+n_{\ell}=t\end{subarray}}\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell}),\\ q_{j}\text{ distinct},\\ q_{j}\nmid N\end{subarray}}$ $\left[\left(\prod_{j=1}^{\ell}\widehat{\phi}\left(\frac{\log q_{j}}{\log Q_{k,N}}\right)^{n_{j}}\frac{\log^{n_{j}}q_{j}}{q_{j}^{n_{j}/2}\log^{n_{j}}Q_{k,N}}\right)\right$.
(3.7) $\left.\times\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j},\\ m_{j}\equiv n_{j}(2)\end{subarray}}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\lambda_{\cdot}\left(\prod\nolimits_{j=1}^{\ell}q_{j}^{m_{j}}\right)\right)\right],$
where all the $c_{n_{j},m_{j}}$ are bounded by $2^{t}$.
@@
(3.8) $\sum_{\begin{subarray}{c}(p_{1},\dots,p_{t})\\ p_{i}\nmid N\end{subarray}}\left[\left(\prod_{i=1}^{t}\widehat{\phi}\left(\frac{\log p_{i}}{\log Q_{k,N}}\right)\frac{\log p_{i}}{\sqrt{p_{i}}\log Q_{k,N}}\right)\ \mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\prod\nolimits_{i=1}^{t}\lambda_{\cdot}(p_{i})\right)\right].$
@@
(3.9) $\frac{1}{\ell!}\binom{t}{n_{1}}\binom{t-n_{1}}{n_{2}}\cdots\binom{n_{\ell}}{n_{\ell}}\ =\ \frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}.$
@@
(3.10) $\left.\times\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(\prod\nolimits_{j=1}^{\ell}\lambda_{\cdot}(q_{j})^{n_{j}}\right)\right].$
As in (2.10), we write $\lambda_{\cdot}(q_{j})^{n_{j}}$ as a linear combination of Hecke eigenvalues at the powers of $q_{j}$:
(3.11) $\lambda_{\cdot}(q_{j})^{n_{j}}\ =\ \sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}(2)\end{subarray}}c_{n_{j},m_{j}}\ \lambda_{\cdot}(q_{j}^{m_{j}}),$
where each $c_{n_{j},m_{j}}$ is bounded by $2^{n_{j}}\leq 2^{t}$ (see (2.12)). Making this substitution in (3.10) gives (3.7). ∎
@@ 4. Weighted Trace Formula
In this section, we derive an asymptotic trace formula for the weighted average of the $m$th Hecke eigenvalue over $\mathcal{F}_{k}(N)^{\prime}$ when $(m,D)=1$:
(4.1) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(\lambda_{\cdot}(m))\ :=\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)\lambda_{f}(m)}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$
Trace formulas are more readily attainable for $\mathcal{F}_{k}(N)$ than $\mathcal{F}_{k}(N)^{\prime}$. This is because the equidistribution laws that govern Hecke eigenvalues—and thereby influence trace formulas—are better understood for the space of all forms $S_{k}(N)$ rather than the space of newforms alone. Studying families of newforms usually entails extensive bookkeeping and adapting trace formulas from the ambient space to the space of newforms, e.g., [ILS00, Propositions 2.1-2.8]. Fortunately, we do not have to do any bookkeeping in our study, owing to the strategic conditioning on $k$ and $N$ in our definitions of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$.
@@
(4.2) $\Lambda\left(\frac{1}{2},g_{N}\times\chi\right)\ =\ \frac{\chi(N)}{N^{k/2}}\Lambda\left(\frac{1}{2},g\times\chi\right)\ =\ 0\implies w_{\chi,r}(g_{N})=0.$
It follows that the weight vanishes for all forms in the span of $g_{N}$, $f$ in particular. This is to say that $w_{\chi,r}(f)$ vanishes for all oldforms $f\in\mathcal{F}_{k}(N)$; hence any $w_{\chi,r}$-weighed average over $\mathcal{F}_{k}(N)$ reduces to a $w_{\chi,r}$-weighted average over $\mathcal{F}_{k}(N)^{\prime}$, and we can refer to these interchangeably.
In the interest of finding a weighted trace formula for (4.1), we call attention to a related formula in [JK15, Theorem 1.1]. For all complex numbers $s=\sigma+it$ in the strip $1-\frac{k-1}{2}<\sigma<\frac{k-1}{2}$, Jackson and Knightly derived a formula for the sum of Fourier coefficients over $\mathcal{F}_{k}(N)$ weighted by
(4.3) $w_{\chi,r}^{\prime}(f)\ :=\ \frac{\Lambda\left(s,f\times\chi\right)\overline{a_{f}(r)}}{\lVert f\rVert^{2}},\quad f\in\mathcal{F}_{k}(N).$
When $k>2$ and $N=1$, for all $m\geq 1$ satisfying $(m,D)=1$, they get
$\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)a_{f}(m)$
$=\$$\ \frac{2^{k-1}(2\pi rm)^{\frac{k-1}{2}-s}}{(k-2)!}\Gamma\left(s+\frac{k-1}{2}\right)\sum_{d\mid(r,m)}d^{2s}\chi\left(\frac{rm}{d^{2}}\right)$
$+\$$\frac{2^{k-1}(2\pi rm)^{\frac{k-3}{2}+s}}{(k-2)!}\Gamma\left(\frac{k+1}{2}-s\right)\frac{i^{k}}{D^{2s-1}}\frac{\tau(\chi)^{2}}{D}\sum_{d\mid(r,m)}d^{-2s+2}\overline{\chi\left(\frac{rm}{d^{2}}\right)}$
[…]
likewise, when $k>2$ and $N>1$, for all $m\geq 1$ with $(m,D)=1$, they get
$\sum_{f\in\mathcal{F}_{k}(N)}w_{\chi,r}^{\prime}(f)a_{f}(m)$
$=\$$\ \frac{2^{k-1}(2\pi rm)^{\frac{k-1}{2}-s}}{(k-2)!}\Gamma\left(s+\frac{k-1}{2}\right)\sum_{d\mid(r,m)}d^{2s}\chi\left(\frac{rm}{d^{2}}\right)$
[…]
The implied constants are provided explicitly in [JK15, (1.5)] and depend only on $s$. As in [KR19, 160], we have adjusted these equations for the fact that [JK15] normalizes the completed $L$-function $\Lambda(s,f\times\chi)$ to have central point $\frac{k}{2}$ rather than $\frac{1}{2}$; that is, we apply the transformation $s\mapsto s+\frac{k-1}{2}$ to [JK15, (1.4)] to satisfy our normalization.
In our application, we take $\chi$ to be a real character, so $\chi=\overline{\chi}$ and $\chi^{2}=1$; and $s=1/2$. We also allow the implied constants to depend on $r$ and $D$ since they are fixed. Specializing (4.4) and (4.5) in this regard, we have that for all $k>2$ and $m\geq 1$ with $(m,D)=1$,
$\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}^{\prime}(f)a_{f}(m)$
$=\$$\frac{2^{k-1}(2\pi)^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+\frac{i^{k}\tau(\chi)^{2}}{D}\right)(rm)^{\frac{k}{2}-1}\chi(rm)\ \sigma_{1}((r,m))$
[…]
$\sum_{f\in\mathcal{F}_{k}(N)}w_{\chi,r}^{\prime}(f)a_{f}(m)$
$=\$$\frac{2^{k-1}(2\pi)^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}(rm)^{\frac{k}{2}-1}\chi(rm)\ \sigma_{1}((r,m))$
[…]
In the following three lemmas, we adapt these asymptotic trace formulas for the weighted average of our interest, i.e., with weights $w_{\chi,r}$ rather than $w_{\chi,r}^{\prime}$: (4.1).
@@
Suppose $k>2$. Let $\sigma_{1}$ be the divisor sum function, and $S_{r}(m):=\sum_{d\mid(r,m)}d\cdot\sigma_{1}\left(\left(r,\frac{rm}{d^{2}}\right)\right)$. Then, for any integer $m\geq 1$ with $(m,D)=1$, the sum of the $m$th Hecke eigenvalue over $\mathcal{F}_{k}(1)$ is equal to
$\sum_{f\in\mathcal{F}_{k}(1)}w_{\chi,r}(f)\lambda_{f}(m)$
$=\$$\frac{2^{k-1}(2\pi r^{2})^{\frac{k}{2}-1}\Gamma\left(\frac{k}{2}\right)}{(k-2)!}\left(1+i^{k}\frac{\tau(\chi)^{2}}{D}\right)m^{-\frac{1}{2}}\chi(m)S_{r}(m)$
[…]
where the implied constant depends only on $r$ and $D$.